How many sheets of drywall are needed for the walls and ceiling of the cabin that is 28 feet x 36 feet. The standard wall height is 8 feet.  ( Hint there are no interior walls) A sheet of drywall is ½” x 4’ x 8’ 


1
Expert's answer
2021-11-16T11:32:05-0500

Area of the walls and ceiling:

"A=(28+36)*2*8+28*36=2032."

Area of the sheet:

"A_s=4*8=32."

Number of sheets:

"N=\frac{A}{a_s}=\frac{2032}{32}=63.5."